OpenStudy (anonymous):

the length of a rectangular banner is 4 feet longer than its width if the area is 96 square feet find the dimension The width of the banner is _____ feet the length is _____ feet

OpenStudy (anonymous):

w(w+4)=96

OpenStudy (anonymous):

w^2+4w=96

OpenStudy (anonymous):

w^2+4w-96=0